OpenStudy (anonymous):

SOMEONE HELP ME PLEEASE!? When solving the system x + 2y = 9 and x - y = 6, which expression could be substituted for x in the first equation? 9 - x y - 9/2 9 - 2y 1/2 y + 9

OpenStudy (mertsj):

The problem says you must substitute for x in the first equation. That means you must get a value for x from the second equation so you have something to substitute. Are you with me so far?

OpenStudy (anonymous):

so find x in the second equation

OpenStudy (mertsj):

That's what your problem says but that is not one of your answer choices so I think you need to repost the problem EXACTLY as it is stated.

OpenStudy (anonymous):

what do you mean this is the prob i copoy and pasted it

OpenStudy (mertsj):

Ok. Solve the first equation for x and choose that answer.

OpenStudy (anonymous):

x=6-y?

OpenStudy (anonymous):

x-y=6 then i did this x=6-y

OpenStudy (mertsj):

Which one is the first equation?

OpenStudy (anonymous):

x + 2y = 9 that is the firts equation right?

OpenStudy (anonymous):

@Mertsj

OpenStudy (uri):

@satellite73 @Hero help her :3

OpenStudy (anonymous):

wheres blondie? :)

OpenStudy (anonymous):

i dont know cant anyone help me? :'(

OpenStudy (anonymous):

x+9-2y and x=6-y

OpenStudy (anonymous):

?

OpenStudy (shaik0124):

how can i help u

OpenStudy (anonymous):

haha that question up there at the very top ^^

OpenStudy (shaik0124):

see u can substitute 3rd option i.e, 9-2y in second euation i.e, x-y=6 what u wrote in the questionis wrong

OpenStudy (shaik0124):

the option u gave is for substituting in second equation not in first

OpenStudy (shaik0124):

shall i show the work

OpenStudy (anonymous):

wait what im a littke confused and yes id like to see how you got it please

OpenStudy (anonymous):

the third one is the same

OpenStudy (shaik0124):